Circular No. 8351

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 Q4 (NEAT) S. M. Brafford, University of Dayton; M. L. Sitko, University of Cincin-

nati; and R. W. Russell and D. L. Kim, The Aerospace Corporation, report 3–13-μm spectrophotometry of comet C/2001 Q4, obtained on May 31.2 UT with the Mt. Lemmon 1.5-m University of Minnesota telescope (+ Aerospace Broadband Array Spectrograph System; 8′′.5 aperture; 49′′ chop throw; integration times 10 min on the comet and 20 min on the reference star, α Lyr): “A smooth comet continuum was seen to rise from 3.5 to 8.4 μm, beyond which a strong silicate emission band was observed. An underlying blackbody continuum with a temperature of ∼ 345 ± 10 K was fit to the continuum flux at 5, 8.4, and 12 μm. This grain temperature is ∼ 24 percent higher than that of an equilibrium blackbody at the heliocentric distance of the comet, and higher than that observed on May 14.1 (IAUC 8342) using the same instrument and telescope when the comet was closer to the sun. Structure is seen in the silicate feature, including emission peaks at 10.5 and 11.2 μm. The silicate-feature-to-continuum ratio was ∼ 1.43 ± 0.04, lower than that reported on May 14.2. No scattered solar flux was detected at the shorter wavelengths after subtracting the thermal continuum. With our aperture, the comet has the following narrowband (∼ 0.25 μm) magnitudes and combined random errors (due to calibration star and comet, as well as variations due to the presence of real spectral structure): [3.7 μm] = 6.8 ± 0.3, [4.7 μm] = 4.60 ± 0.11, [5 μm] = 4.27 ± 0.06, [8 μm] = 1.71 ± 0.06, [10.5 μm] = –0.54 ± 0.02, and [12 μm] = –0.49 ± 0.03 (the stated errors are standard deviations of the mean).”

COMETS C/2003 V2, V3, V4, V5 (SOHO) Further to IAUC 8348, K. Battams reports measurements for addi-

tional Kreutz sungrazing comets found on SOHO website C2 images (by R. Kracht, C/2003 V2; M. Boschat, C/2003 V3, V4, and V5); C/2003 V3 and V5 were also visible on C3 images. The reductions by R. Kisala and orbital elements by B. G. Marsden appear on MPEC 2004-K34.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 V2 Nov. 1.896  14^{h}21.^{m}9  −15^{o}35′
C/2003 V3  3.513  14 19.5  −17 14
C/2003 V4  4.938  14 33.7  −16 43
C/2003 V5  5.388  14 23.2  −20 18
2004 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green